# The radiation condition for Helmholtz equations above locally perturbed periodic surfaces

### v1

@@ Abstract
The radiation condition is the key question in the mathematical modelling for scattering problems in unbounded domains. Mathematically, it plays the role as the ”boundary condition” at the infinity, which guarantees the well-posedness of the mathematical problem; physically, it describes the behaviour of the physical waves. In this paper, we focus on the radiation conditions for scattering problems with periodic media embedded in two dimensional half-spaces. According to Hu et al. (2021), the radiating solution satisfies the Sommerfeld radiation condition:
@@
Although there are literature which have studied this problem, there is no specific method for dealing with periodic structures. Due to this reason, the important properties for the periodic structures are always ignored. Moreover, the existing method is not extendable to bi-periodic structures in three dimensional spaces.
In this paper, we study the radiation condition for the scattering problem with periodic medium, which is modelled by the Helmholtz equation. We introduce a novel method based on the Floquet-Bloch transform, which, to the best of the author’s knowledge, is the first method that works particularly for periodic media. With this method, we improve the Sommerfeld radiation condition for the scattered field from periodic media to:
@@
Moreover, the prospect of extending this method to 3D cases is optimistic.
@@ 1 Introduction
For the mathematical modelling and analysis of the scattering problems with unbounded domains, the radiation condition has been the central and key topic. From the mathematical point of view, we need a ”boundary condition” at the infinity, to formulate a well-posed mathemtaical problem to describe the physical process. From the physical point of view, a proper radiation condition describes the physical solution explicitly, thus the mathematical model coincides with the physical phenomenan in the real world. For the easiest case, i.e., time-harmonic acoustic waves scattered by a bounded obstacle, the scattered field satisfies the well-known Sommerfeld radiation condition (see [33]). Similarly, for electromagnetic waves the radiation condition is know as the Silver-Müller condition. For details we refer to [13].
@@
In this paper, we consider a much more complex case, i.e., time-harmonic acoustic waves scattered by periodic media. The scattering problems in periodic structures have a large area of applications in, for example, photonic crystals. We refer to [32] for details. The problem is modelled by the Helmholtz equation, where the domain is a periodic half-space. Compared to the bounded obstacles, this problem is more difficult due to the unbounded domains. If we ignore the periodicity of the structure but simply treat it as a infinite layer or interface, then the theory of rough surface/layer scattering problems can be adopted. Roughly speaking, the theory is based on the integral equation representation for the solutions, and detailed analysis on the integrals in infinite curves is carried out to study the properties of the radiation conditions. We refer to [2, 7, 8, 9, 10, 11] for a series of earlier work related to this topic. For details we refer to [18]. Here we also mention that, for the step-like surfaces, a shape condition is obtained, see [28]. Based on the method, Hu, Lu and Rathsfeld studied the radiation condition for the Dirichlet periodic surface scattering problems. They proved that the scattered field satisfies the half-plane Sommerfeld radiation condition (HPSRC) in a half space above the periodic surface:
@@
$\sup_{x_{2}>h}r^{1/2}\left|\frac{\partial u}{\partial r}(x)-\mathrm{i}ku(x)\right|\rightarrow 0,\quad\sup_{x_{2}>h}r^{1/2}\left|u(x)\right|<\infty,\quad r\rightarrow\infty.$
@@
The general case, i.e., scattering problems with periodic inhomogeneous layers, it is more complex due to the existence of guided waves (see [4]). Based on his previous work as well as the work in [18], Kirsch proved the radiation condition for the radiating part in [20]. At the end of that paper, he also gave the asymptotic behaviour of the radiating part. However, due to technical difficulties, the extension of this kind of problems to bi-periodic structures in three dimensional spaces is not realistic.
@@
We can clearly observe that, in all the mentioned work, the analysis is based on the rough surface/layer scattering problem, instead of the periodic surface/layer scattering problems. It implies that some major properties for the periodic structures may be ignored in the analysis, and the results are expected to be improved if specific methods are developed for the periodic structures. Motivated by this, we intend to introduce a novel method, which is, to the best of the author’s knowledge, the first to treat particularly the radiation conditions for the periodic structures. We will utilize the properties for the periodic problems and obtain a shaper radiation condition compared to HPSRC introduced by Hu, Lu and Rathsfeld:
@@
$\sup_{x_{2}>h}r^{3/2}\left|\frac{\partial u}{\partial r}(x)-\mathrm{i}ku(x)\right|<\infty,\quad\sup_{x_{2}>h}r^{1/2}\left|u(x)\right|<\infty,\quad r\rightarrow\infty.$
@@
An important tool to treat periodic structures is the so-called Floquet-Bloch transform. We refer to [17, 15, 22, 35] for the analysis of the radiation conditions for scattering problems in closed periodic waveguides. For scattering problems with (locally perturbed) periodic surfaces, we refer to [12, 25, 16, 24, 26, 27, 34] for the applications in numerical simulations. For more general open waveguide problems, i.e., with the existence of guided waves, we refer to [21] for detailed discussions. In this paper, we will also apply the Floquet-Bloch transform to the original scattering problem. The original problem is then written into a equivalent family of quasi-periodic problems, and the solution has the form of the integral of the quasi-periodic functions with respect to the quasi-periodicity parameter. With the integral representation, the scattered field is decomposed into an evanescent wave and a Herglotz wave function with particular density function. Thus we only need to investigate the radiation conditions for both of the functions. Technically, the asymptotic behaviours of some parameter dependent integrals will be required for the investigation.
The rest of the paper will be organized as follow. In Section 2, the mathematical model for the problem will be given and we will also give a brief review of some important notations and results. The Floquet-Bloch transform and its applications will be described in Section 3. For the convenience of the readers, we will show the main theorem and also the sketch of the proof in Section 4. Afterwards, we first introduce the asymptotic behaviours of some special parameter dependent integrals in Section 5, and the decay of the functions in a bounded strip is given in Section 6. For the estimation of radiation conditions with three different form of waves, we refer to Sections 7, 8 and 9.
@@ 2 The Dirichlet scattering problems
Let $\Gamma^{h}:=\mathbb{R}\times\{h\}$ for any $h\in\mathbb{R}$, then $\Gamma^{H}$ is the straight line which lies above $\Gamma$. Then $\Omega^{H}$ is defined as the domain between $\Gamma$ and $\Gamma^{H}$. Thus $\Omega^{H}$ is an infinite strip which is $2\pi$-periodic in $x_{1}$-direction. Let $\mathbb{R}^{+}_{h}$ be the half space above $\Gamma^{h}$, i.e., $\mathbb{R}^{+}_{h}:=\mathbb{R}\times(h,\infty)$. For the visulization we refer to Figure 1.
@@
where $y$ is the source located above $\Gamma$; $G(x,y)$ is the half-space Green’s function given by
@@
here $H_{0}^{(1)}(\cdot)$ is the Hankel function of the first kind, $y^{\prime}$ is the reflection of $y$ with respect to $x_{2}=0$. Moreover, it is know that for the radiating solution $u$ satisfies the so-called upward propagationg radiation condition (UPRC, see [9]):
@@
The UPRC is equivalent to the following DtN map in terms of the Fourier transform (see [6]):
@@
$H_{s}^{r}(D):=\left\{\phi:\,(1+|x|^{2})^{r/2}\phi(x)\in H^{s}(D)\right\}.$
Note that this theorem works not only for periodic surfaces, but also for more general cases: the graph of bounded functions.
@@
For the problem we are discussing, when $f$ is given by $-G(x,y)$ for a fixed $y$ above $\Gamma$, it is already known that $f$ decays at the rate of $O\left(|x_{1}|^{-3/2}\right)$ (see [7]). Thus it is easily checked that $f\in H_{r}^{1/2}$ for any $r<1$. Thus $u\in H_{r}^{1}(\Omega^{H})$.
Although we only consider the incident field $-G(x,y)$, the result is easily extendable to other cases, e.g., $\Phi(x,y)$, or the Herglotz wave functions.
The Dirichlet boundary condition here is not changable. If we choose another boundary condition, the uniqueness result in Theorem 2.1 may fail. Similarly, we always need to restrict $\Gamma$ to be a graph of a function.
@@ 3.1 The Floquet-Bloch transform, and definitions of spaces
The transform is well-defined for any $\phi\in C_{0}^{\infty}(\mathbb{R})$ since there are only finite number of non-vanishing terms in the series. But the operator can also be extended to more general function spaces. For the explicit definition of the space $H_{0}^{r}\left(\Lambda;H^{s}_{\alpha}[-\pi,\pi]\right)$ we also refer to [24]. Then we have the following properties of the operator $\mathcal{J}$.
@@
$(\mathcal{J}^{-1}\psi)(x+2\pi j)=\int_{-1/2}^{1/2}\psi(\alpha,x)e^{\mathrm{i}2\pi j\alpha}\,\mathrm{d}\alpha\,,\quad x\in[-\pi,\pi],\,j\in\mathbb{Z}.$ (3.2)
$\Omega^{H}$$\Gamma$$\Gamma^{H}$$\mathcal{J}$${\mathcal{J}}^{-1}$$\Omega^{H}_{0}$$\Gamma_{0}$$\Gamma^{H}_{0}$$\Lambda$
@@
Note that the definition is easily extend to the two-dimensional domains, which are periodic in the $x_{1}$-dimension. For convenience, we first define the restrictions of the domains and curves in one periodicity cell:
@@
$C^{\infty}(K;S(W)):=$$\left\{f\in\mathcal{D}^{\prime}(K\times W):\,\text{ for all }j=0,1,\dots,\infty,\,\frac{\partial^{j}f(\alpha,\cdot)}{\partial\alpha^{j}}\in S(W),\right.$
[…]
We also define the spaces $C^{n}_{0}(K;S(W))$ and $C^{\infty}_{0}(K;S(W))$ as subspaces of $C^{n}(K;S(W))$ and $C^{\infty}(K;S(W))$, where the functions are compactly supported in $K$ with respect to $\alpha$; and $C^{n}_{\mathrm{per}}(K;S(W))$ and $C^{\infty}_{\mathrm{per}}(K;S(W))$ as subspaces of $C^{n}(K;S(W))$ and $C^{\infty}(K;S(W))$, where the functions are periodic.
@@ 3.2 Application of the Floquet-Bloch transform and results
In this section, we apply the Floquet-Bloch transform to the unique solution $u$ to the problem (2.1). First, Theorem 2.1 and Theorem 3.2 guarantees that the transform works. Let $w(\alpha,x):=\mathcal{J}u$, then for each fixed $\alpha\in\Lambda$, $w(\alpha,\cdot)$ is $\alpha$-quasi-periodic and satisfies
@@
To prepare for the proof for the main results, the existence and regularity of the transformed field $w$ as described in Corollary 3.4 is not sufficient. We need more detailed analysis for the field $w$. We refer to [34] for more explicit descriptions for the regularity with respect to $\alpha$. The results are summarized as follows.
@@
where $w_{0},\,w_{1},\,w_{2}\in C^{\infty}\left(\Lambda;H^{1}_{\alpha}(\Omega^{H}_{0})\right)$. Moreover, $w_{0}\in C^{\infty}_{\mathrm{per}}\left(\Lambda;H^{1}_{\alpha}(\Omega^{H}_{0})\right)$ and $w_{1},\,w_{2}\in C^{\infty}_{0}\left((\pm\alpha_{0}-\delta,\pm\alpha_{0}+\delta);H^{1}_{\alpha}(\Omega^{H}_{0})\right)$. Moreover, the interval $\Lambda$ is adjusted such that $[-\alpha_{0}-\delta,-\alpha_{0}+\delta]$ and $[\delta_{0}-\delta,\alpha_{0}+\delta]$ are both in the interior of $\overline{\Lambda}$.
@@
where $w_{0},\,w_{1},\,w_{2}\in C^{\infty}\left(\Lambda;H^{1}_{\alpha}(\Omega^{H}_{0})\right)$. Moreover, $w_{0}\in C^{\infty}_{\mathrm{per}}\left(\Lambda;H^{1}_{\alpha}(\Omega^{H}_{0})\right)$ and $w_{1},\,w_{2}\in C^{\infty}_{0}\left((\alpha_{0}-\delta,\pm\alpha_{0}+\delta);H^{1}_{\alpha}(\Omega^{H}_{0})\right)$. Also, the parameter $\delta>0$ is chosen such that $[\alpha_{0}-\delta,\alpha_{0}+\delta]$ lies in the interior of $\overline{\Lambda}$.
From now on, define the set

### v2

# The radiation condition for Helmholtz equations above (locally perturbed) periodic surfaces

@@ Abstract
The radiation condition is the key question in the mathematical modelling for scattering problems in unbounded domains. Mathematically, it plays the role as the ”boundary condition” at the infinity, which guarantees the well-posedness of the mathematical problem; physically, it describes the far-field asymptotic behaviour of the physical waves. In this paper, we focus on the radiation conditions for scattering problems above (locally perturbed) periodic surfaces. According to Hu et al. (2021), the radiating solution satisfies the Sommerfeld radiation condition:
@@
Although there are literature which have studied this problem, there is no specific method for dealing with periodic structures. Due to this reason, the important properties for the periodic structures may be ignored. Moreover, the existing method is not extendable to bi-periodic structures in three dimensional spaces.
In this paper, we study the radiation condition for the time-harmonic scattering problem with periodic surfaces, which is modelled by the Helmholtz equation. We introduce a novel method based on the Floquet-Bloch transform, which, to the best of the author’s knowledge, is the first method that works particularly for periodic media. With this method, we improve the Sommerfeld radiation condition for the scattered field from periodic media to:
@@
More importantly, the prospect of extending this method to 3D cases is optimistic.
@@ 1 Introduction
For the mathematical modelling and analysis of the scattering problems with unbounded domains, the radiation condition is always the key topic. From the mathematical point of view, we need a ”boundary condition” at the infinity, to formulate a well-posed mathemtaical problem to describe the physical process. From the physical point of view, a proper radiation condition describes the physical solution explicitly, such that the mathematical model coincides with the physical phenomenan in the real world. For the easiest case, i.e., time-harmonic acoustic waves scattered by a bounded obstacle, the scattered field satisfies the well-known Sommerfeld radiation condition (see [30]). Similarly, for electromagnetic waves the radiation condition is know as the Silver-Müller condition. For details we refer to [12].
@@
In this paper, we consider a much more complex case, i.e., time-harmonic acoustic waves scattered by (locally perturbed) periodic surfaces. The scattering problems in periodic structures have a large area of applications in, for example, photonic crystals. We refer to [29] for details. The problem is modelled by the Helmholtz equation, where the domain is a periodic half-space. Compared to the bounded obstacles, this problem is more difficult due to the unbounded domains and complex singularities. If we ignore the periodicity of the structure but simply treat it as a infinite layer or interface, then the theory of rough surface/layer scattering problems can be adopted. Roughly speaking, the theory is based on the integral equation representation for the solutions, and detailed analysis on the integrals in infinite curves is carried out to study the properties of the radiation conditions. We refer to [2, 6, 7, 8, 9, 10] for a series of earlier work related to this topic. Here we also mention that, for the special step-like surfaces, a shaper condition is obtained, see [26]. Based on the method, Hu, Lu and Rathsfeld studied the radiation condition for the Dirichlet periodic surface scattering problems. They proved that the scattered field satisfies the half-plane Sommerfeld radiation condition (HPSRC) in a half space above the periodic surface:
@@
$\sup_{x_{2}>h}r^{1/2}\left|\frac{\partial u}{\partial r}(x)-\mathrm{i}ku(x)\right|\rightarrow 0,\quad\sup_{x_{2}>h}r^{1/2}\left|u(x)\right|<\infty,\quad|x|=r\rightarrow\infty.$
@@
For details we refer to [17]. The general case, i.e., scattering problems with periodic inhomogeneous layers, it is more complex due to the existence of guided waves (see [3]). Based on his previous work as well as the work in [17], Kirsch proved the radiation condition for the radiating part in [19]. At the end of that paper, he also gave the asymptotic behaviour of the radiating part. However, due to technical difficulties, the extension of this kind of problems to bi-periodic structures in three dimensional spaces is not realistic.
@@
We can clearly observe that, in all the mentioned work, the analysis is based on the rough surface/layer scattering problem, instead of the periodic surface/layer scattering problems. It implies that some major properties of the periodic structures may be ignored in the analysis, and the results are expected to be improved if specific methods are developed. Motivated by this, we intend to introduce a novel method, which is, to the best of the author’s knowledge, the first one to treat particularly the radiation conditions for the periodic structures. We will utilize the properties for the periodic problems and obtain a shaper radiation condition compared to HPSRC introduced by Hu, Lu and Rathsfeld in [17]:
@@
$\sup_{x_{2}>h}r^{3/2}\left|\frac{\partial u}{\partial r}(x)-\mathrm{i}ku(x)\right|<\infty,\quad\sup_{x_{2}>h}r^{1/2}\left|u(x)\right|<\infty,\quad|x|=r\rightarrow\infty.$
@@
An important tool to treat periodic structures is the so-called Floquet-Bloch transform. We refer to [16, 14, 21, 32] for the analysis of the radiation conditions for scattering problems in closed periodic waveguides. For scattering problems with (locally perturbed) periodic surfaces, we refer to [11, 23, 15, 22, 24, 25, 31] for the applications in numerical simulations. For more general open waveguide problems, i.e., with the existence of guided waves, we refer to [20] for detailed discussions. In this paper, we will also apply the Floquet-Bloch transform to the scattering problem. The original problem is then written into a equivalent family of quasi-periodic problems, and the solution has the form of the integral of the quasi-periodic functions with respect to the quasi-periodicity parameter. With the integral representation, the scattered field is decomposed into an evanescent wave and a Herglotz wave function with particular density function. Thus we only need to investigate the radiation conditions for both of the functions. Technically, the asymptotic behaviours of some parameter dependent integrals will be required for the investigation.
The rest of the paper will be organized as follow. In Section 2, the mathematical model for the problem with purely periodic surface will be given and we will also give a brief review of some important notations and results. The Floquet-Bloch transform and its applications will be described in Section 3. For the convenience of the readers, we will present the main theorem (Theorem 4.2) and also the sketch of the proof in Section 4. The proof of Part (A) in Theorem 4.2 we refer to Section 5. The proof of Part (B) is much more challenging. We prove three sub-topics in Sections 6, 7 and 8, and then summarize the proof in Section 9. Finally, we extend the proof for locally perturbed periodic surfaces in Section 10.
@@ 2 The Dirichlet scattering problems
Let $\Gamma^{h}:=\mathbb{R}\times\{h\}$ for any $h\in\mathbb{R}$, then $\Gamma^{H}$ is the straight line which lies above $\Gamma$. Let $\Omega^{H}$ be defined as the domain between $\Gamma$ and $\Gamma^{H}$. Thus $\Omega^{H}$ is an infinite strip which is $2\pi$-periodic in $x_{1}$-direction. Let $U_{h}$ be the half space above $\Gamma^{h}$, i.e., $U_{h}:=\mathbb{R}\times(h,\infty)$. For the visulization we refer to Figure 1.
@@
where $y=(y_{1},y_{2})$ is the source located above $\Gamma$ (i.e., $y_{2}>\zeta(y_{1})$); $G(x,y)$ is the half-space Green’s function given by
@@
here $H_{0}^{(1)}(\cdot)$ is the Hankel function of the first kind, $y^{\prime}=(y_{1},-y_{2})$ is the reflection of $y$ with respect to $x_{2}=0$. Moreover, it is know that for the radiating solution $u$ satisfies the so-called upward propagationg radiation condition (UPRC, see [8]):
@@
The UPRC is equivalent to the angular spectrum representation (see [2]),
$u(x_{1},x_{2})=\int_{\Gamma^{H}}e^{\mathrm{i}\xi x_{1}+\mathrm{i}\sqrt{k^{2}-\xi^{2}}(x_{2}-H)}\hat{u}(\xi,H)\,\mathrm{d}\xi\,,$ (2.3)
which is equivalent to the DtN map in terms of the Fourier transform (see [5]):
@@
$H_{s}^{r}(D):=\left\{\phi:\,(1+|x|^{2})^{r/2}\phi(x)\in H^{s}(D)\right\},$
where $H^{s}(D)$ is the classic Sobolev space. Note that this theorem works not only for periodic surfaces, but also for more general rough surfaces with Dirichlet boundary conditions.
@@
For the problem we are discussing, when $f$ is given by $-G(x,y)$ for a fixed $y$ above $\Gamma$, it is already known that $f$ decays at the rate of $O\left(|x_{1}|^{-3/2}\right)$ (see [6]). Thus it is easily checked that $f\in H_{r}^{1/2}$ for any $r<1$. Thus $u\in H_{r}^{1}(\Omega^{H})$ is the unique solution.
The Dirichlet boundary condition here is not changable. If we choose another boundary condition, the uniqueness result in Theorem 2.1 may fail. Moreover, we always need to restrict $\Gamma$ to be a graph of a function.
@@ 3.1 The Floquet-Bloch transform, and definitions of spaces
The transform is well-defined for any $\phi\in C_{0}^{\infty}(\mathbb{R})$ since there are only finite number of non-vanishing terms in the series. But the operator can also be extended to more general function spaces. We give two examples of the norm of the space $H_{0}^{r}\left(\Lambda;H^{s}_{\alpha}[-\pi,\pi]\right)$ when $s=r=0$ (in this case, it is also denoted by $L^{2}\left(\Lambda;L^{2}[-\pi,\pi]\right)$) and $s\in\mathbb{R}$ and $r=n\in\mathbb{N}$:
$\|\psi\|_{L^{2}\left(\Lambda;L^{2}[-\pi,\pi]\right)}$ $=\left[\int_{\Lambda}\int_{-\pi}^{\pi}|\psi(\alpha,x)|^{2}\,\mathrm{d}x\,\,\mathrm{d}\alpha\,\right]^{1/2};$
$\|\psi\|_{H_{0}^{n}\left(\Lambda;H^{s}[-\pi,\pi]\right)}$ $=\left[\sum_{\ell=0}^{n}\int_{\Lambda}\left\|\frac{\partial^{\ell}\psi}{\partial\alpha^{\ell}}(\alpha,\cdot)\right\|^{2}_{H^{s}_{\alpha}(-\pi,\pi)}\,\mathrm{d}\alpha\,\right]^{1/2}.$
For the explicit definition of the space $H_{0}^{r}\left(\Lambda;H^{s}_{\alpha}[-\pi,\pi]\right)$ for general $r,s\in\mathbb{R}$, we also refer to [22]. Then we have the following properties of the operator $\mathcal{J}$.
@@
$(\mathcal{J}^{-1}\psi)(x+(2\pi,0)^{\top})=\int_{-1/2}^{1/2}\psi(\alpha,x)e^{\mathrm{i}2\pi j\alpha}\,\mathrm{d}\alpha\,,\quad x\in[-\pi,\pi],\,j\in\mathbb{Z}.$ (3.2)
$\Omega^{H}$$\Gamma$$\Gamma^{H}$$\mathcal{J}$${\mathcal{J}}^{-1}$$x\in\Omega^{H}_{0}$$\Gamma_{0}$$\Gamma^{H}_{0}$$\alpha\in\Lambda$
@@
Note that the definition is easily extended to the two-dimensional domains, which are periodic in the $x_{1}$-dimension. For convenience, we first define the restrictions of the domains and curves in one periodicity cell:
@@
$C^{\infty}(K;S(W)):=\bigcap_{n=1}^{\infty}C^{n}(K;S(W)).$
We also define the spaces $C^{n}_{0}(K;S(W))$ and $C^{\infty}_{0}(K;S(W))$ as subspaces of $C^{n}(K;S(W))$ and $C^{\infty}(K;S(W))$, where the functions are compactly supported in $K$ with respect to $\alpha$; and $C^{n}_{\mathrm{per}}(K;S(W))$ and $C^{\infty}_{\mathrm{per}}(K;S(W))$ as subspaces of $C^{n}(K;S(W))$ and $C^{\infty}(K;S(W))$, where the functions are periodic with respect to $\alpha$.
@@ 3.2 Application of the Floquet-Bloch transform and results
In this section, we apply the Floquet-Bloch transform to the unique solution $u$ to the problem (2.1). From Theorem 2.1, the solution $u\in H_{r}^{1}(\Omega)$. Then with Theorem 3.2, the Floquet-Bloch tranform is well defined for $u$. Let $w(\alpha,x):=\mathcal{J}u$, then $w\in H_{0}^{r}\left(\Lambda;H^{1}_{\alpha}(\Omega^{H}_{0})\right)$. For each fixed $\alpha\in\Lambda$, $w(\alpha,\cdot)$ is $\alpha$-quasi-periodic and satisfies
@@
To prove the main results, the existence and regularity of the transformed field $w$ as described in Corollary 3.4 is not sufficient. We need more detailed analysis for the field $w$. We refer to [31] for more explicit descriptions for the regularity with respect to $\alpha$. The results are summarized as the following theorem.
Theorem 3.5.
Suppose $\Gamma\subset\mathbb{R}^{2}$ is the graph of a periodic and bounded function $\zeta$, $u$ is the solution to (2.1)-(2.2). Let $w$ be the Floquet-Bloch transform of $u$, then the regularity of $w$ depending on $\alpha$ are discussed with two different cases.
@@
where $w_{0},\,w_{1},\,w_{2}\in C^{\infty}\left(\Lambda;H^{1}_{\alpha}(\Omega^{H}_{0})\right)$. Moreover, $w_{0}\in C^{\infty}_{\mathrm{per}}\left(\Lambda;H^{1}_{\alpha}(\Omega^{H}_{0})\right)$ and $w_{1},\,w_{2}\in C^{\infty}_{0}\left((\pm\alpha_{0}-\delta,\pm\alpha_{0}+\delta);H^{1}_{\alpha}(\Omega^{H}_{0})\right)$. Moreover, the parameter $\delta>0$ is adjusted such that $[-\alpha_{0}-\delta,-\alpha_{0}+\delta]$ and $[\delta_{0}-\delta,\alpha_{0}+\delta]$ are both in the interior of $\overline{\Lambda}$. In this case,
$u(x)=\int_{\Lambda}w_{0}(\alpha,x)\,\mathrm{d}\alpha\,+\int_{\alpha_{0}-\delta}^{\alpha_{0}+\delta}\sqrt{\alpha-\alpha_{0}}\,w_{1}(\alpha,x)\,\mathrm{d}\alpha\,+\int_{-\alpha_{0}-\delta}^{-\alpha_{0}+\delta}\sqrt{\alpha+\alpha_{0}}\,w_{2}(\alpha,x)\,\mathrm{d}\alpha\,,\quad x\in\Omega^{H},$ (3.8)
where $w_{0}$, $w_{1}$ and $w_{2}$ are extended into $\alpha$-quasi-periodic functions for $x\in\Omega^{H}$.
@@
where $w_{0},\,w_{1},\,w_{2}\in C^{\infty}\left(\Lambda;H^{1}_{\alpha}(\Omega^{H}_{0})\right)$. Moreover, $w_{0}\in C^{\infty}_{\mathrm{per}}\left(\Lambda;H^{1}_{\alpha}(\Omega^{H}_{0})\right)$ and $w_{1},\,w_{2}\in C^{\infty}_{0}\left((\alpha_{0}-\delta,\pm\alpha_{0}+\delta);H^{1}_{\alpha}(\Omega^{H}_{0})\right)$. Also, the parameter $\delta>0$ is chosen such that $[\alpha_{0}-\delta,\alpha_{0}+\delta]$ lies in the interior of $\overline{\Lambda}$. In this case,
$u(x)=\int_{\Lambda}w_{0}(\alpha,x)\,\mathrm{d}\alpha\,+\int_{\alpha_{0}-\delta}^{\alpha_{0}+\delta}\sqrt{\alpha-\alpha_{0}}\,w_{1}(\alpha,x)\,\mathrm{d}\alpha\,+\int_{\alpha_{0}-\delta}^{\alpha_{0}+\delta}|\alpha-\alpha_{0}|\,w_{2}(\alpha,x)\,\mathrm{d}\alpha\,,\quad x\in\Omega^{H},$ (3.10)
where again, $w_{0}$, $w_{1}$ and $w_{2}$ are extended into $\alpha$-quasi-periodic functions for $x\in\Omega^{H}$.
From now on, define the set $\mathcal{S}$ and interval $\Lambda$ for different $k$: